Substituting x = – 3 in
 x^2 + 5x + 4 gives (– 3)^2 +5(– 3)+4 = 9+2+4 = 15

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Substituting x =–3 in x^2+5x+4, we have
x^2+5x+4 = (–3)^2+5(–3)+4 = 9–15+4 = - 2.
This is the correct answer.

Thanks